OpenStudy (anonymous):

sqrt{x+2} = 6 - sqrt{7x+2}

OpenStudy (anonymous):

\[\sqrt{x+2} = 6 - \sqrt{7x+2}\]

jhonyy9 (jhonyy9):

square bothe sides (x+2)=36-12sqrt(7x+2)+(7x+2) x+2-36-7x-2=-12sqrt(7x+2) -36-6x=-12sqrt(7x+2) /:(-6) 6+x=2sqrt(7x+2) square both sides 36+12x+x2=4(7x+2) 36+12x+x2 -28x-8=0 x2 -16x +28 =0